OpenStudy (anonymous):

sin(180-theta)sin(180+theta)cosec^2theta is equal to

OpenStudy (anonymous):

sin (a+b) = sina cos b + sinb cos a sin (a -b) = sina cos b - sinb cos a cosc a=\(\dfrac{1}{sin a}\) if your calculating was right, the answer will be -1

OpenStudy (anonymous):

sin180 costheta+sintheta cos180

OpenStudy (anonymous):

sin180 cos theta-sin theta cos180

OpenStudy (anonymous):

@ooops

OpenStudy (anonymous):

The quadratic equation whose roots are sin 18° and cos 36° is

OpenStudy (anonymous):

for the first one sin180 =0, cos 180 =-1

OpenStudy (anonymous):

for the second question, it is (x-sin18)(x-cos 36)=0, open the bracket, you have the required quadratic.

OpenStudy (anonymous):

x^2-xcos36-xsin18-sin18cos36

OpenStudy (anonymous):

you can let it as it is or find out the value of sin 18 and cos 36 from the calculator. I don't have calculator with me, so I can't check.

OpenStudy (anonymous):

only by using calculator u can check it.otherwise no formula